# PSDNorm: Test-Time Temporal Normalization for Deep Learning in Sleep Staging

### v2

@@ Abstract
Distribution shift poses a significant challenge in machine learning, particularly in biomedical applications using data collected across different subjects, institutions, and recording devices, such as sleep data. While existing normalization layers, BatchNorm, LayerNorm and InstanceNorm, help mitigate distribution shifts, when applied over the time dimension they ignore the dependencies and auto-correlation inherent to the vector coefficients they normalize. In this paper, we propose PSDNorm that leverages Monge mapping and temporal context to normalize feature maps in deep learning models for signals. Notably, the proposed method operates as a test-time domain adaptation technique, addressing distribution shifts without additional training. Evaluations with architectures based on U-Net or transformer backbones trained on 10K subjects across 10 datasets, show that PSDNorm achieves state-of-the-art performance on unseen left-out datasets while being 4-times more data-efficient than BatchNorm.
@@ 1 Introduction
Machine learning techniques have achieved remarkable success in various domains, including computer vision, biology, audio processing, and language understanding. However, these methods face significant challenges when there are distribution shifts between training and evaluation datasets [1]. For example, in biological data, such as electroencephalography (EEG) signals, the distribution of the data can vary significantly. Indeed, data is collected from different subjects, electrode positions, and recording conditions. This paper focuses on sleep staging, a clinical task that consists in classifying periods of sleep in different stages based on EEG signals [2]. Depending on the dataset, the cohort can be composed of different age groups, sex repartition, health conditions, and recording conditions [3, 4, 5]. Such variability brings shift in the distribution making it challenging for the model to generalize to unseen datasets.
@@
Normalization layers are widely used in deep learning to improve training stability and generalization. Common layers include BatchNorm [6], LayerNorm [7], and InstanceNorm [8], which respectively compute statistics across the batch, normalize across all features within each sample, and normalize each channel independently within a sample. Some normalization methods target specific tasks, such as EEG covariance matrices [9] or time-series forecasting [10], but they do not fully address spectral distribution shifts reflected in the temporal auto-correlations of signals. In sleep staging, a simple normalization is often applied as preprocessing, e.g., standardizing signals over entire nights [11] or short temporal windows [12]. Recent studies [13, 14] highlight the importance of considering temporal correlation and spectral content in normalization, proposing Temporal Monge Alignment (TMA), which aligns Power Spectral Density (PSD) to a common reference using Monge mapping, going beyond simple z-score normalization. However, these methods remain preprocessing steps that cannot be inserted as layers in the network architecture as it is done with BatchNorm, LayerNorm or InstanceNorm.
Domain Generalization
@@
Sleep staging has been addressed by various neural network architectures, which process raw signals [12, 15, 16], spectrograms [17, 18], or both [19]. More recent approaches involve transformer-based models that handle multimodal [20], spectrogram [21], or heterogeneous inputs [22], offering improved modeling of temporal dependencies. However, most existing models are trained on relatively small cohorts, typically consisting of only a few hundred subjects, which limits their ability to generalize to diverse clinical settings. Notable exceptions include U-Sleep [15], which was trained on a large-scale dataset and incorporates BatchNorm layers to mitigate data variability, and foundational models [23, 24, 25] that achieve strong generalization from vast amount of data but require significant computational resources and are challenging to adapt without fine-tuning. Our focus is on developing smaller, efficient models that balance good generalization with ease of training and deployment in clinical practice.
Test-time Domain Adaptation
Domain generalization involves training models on data from multiple domains without using domain labels, with the goal of achieving broad generalization. In contrast, domain adaptation (DA) utilizes domain information to improve performance by aligning source and target distributions during training [26, 27]. While DA methods have been applied to address domain shifts in sleep staging [28], they typically require access to source data for each new target, which can be limiting. Test-time DA offers a more flexible approach, adapting the model during inference without requiring access to source data [13]. This makes test-time DA particularly well-suited for clinical applications where retraining or data sharing is impractical.
@@
Figure 1: Description of normalization layers. The input shape is $(N,c,\ell)$ with batch size $N$, channels $c$, and signal length $\ell$. BatchNorm estimates the mean $\widehat{\mu}$ and variance $\widehat{\sigma}^{2}$ over batch and time, and learns parameters $(\gamma,\beta)$ to normalize the input. PSDNorm estimates PSDs $\widehat{\mathbf{P}}$ over time and accounts for local temporal correlations. It computes the barycenter PSD $\widehat{\overline{\mathbf{P}}}$, updates it via a running Riemannian barycenter (6), and applies the filter $\widehat{\mathbf{H}}$ to normalize the input. The hyperparameter $F$ controls the extent of temporal correlation considered, thereby adjusting the strength of the normalization.
@@
In this work, we introduce the PSDNorm deep learning layer, a novel approach to address distribution shifts in machine learning for signals. PSDNorm leverages Monge Mapping to incorporate temporal context and normalize feature maps effectively. As a test-time DA method, it adapts to distribution shifts during inference without extra training or source domain access. Compared to a normalization like LayerNorm or InstanceNorm, PSDNorm exploits the sequential nature of the intermediate layer representations as illustrated in Figure 1. We evaluate PSDNorm through extensive experiments on 10 sleep datasets. This evaluation covers 10M of samples across 10K subjects, using a leave-one-dataset-out (LODO) protocol with 3 different random seeds. To the best of our knowledge, such a large-scale and systematic evaluation has never been conducted before. PSDNorm achieves state-of-the-art performance and requires 4 times fewer labeled data to match the accuracy of the best baseline. Results highlight the potential of PSDNorm as a practical and efficient solution for tackling domain shifts in signals.
The paper is structured as follows: Section 2 discusses existing normalization layers and pre-processing. Section 3 introduces PSDNorm, followed by numerical results in Section 4.
@@ 2 Related Works
In this section, we first review fundamental concepts of normalization layers. Then, we recall the Temporal Monge Alignment (TMA) method [13] that aligns the PSD of signals using optimal transport.
@@
Normalization layers enhance training and robustness in deep neural networks. The most common are BatchNorm [6], InstanceNorm [8], and LayerNorm [7]. BatchNorm normalizes feature maps using batch and time statistics, ensuring zero mean and unit variance. The output is adjusted with learnable parameters. InstanceNorm normalizes each channel per sample using its own statistics, independent of the batch (see Fig. 1). Popular in time-series forecasting, it is used in RevIN [10], which reverses normalization after decoding. LayerNorm normalizes across all channels and time steps within each sample, with learnable scaling and shifting. While these normalization layers are widely employed, they operate on vectors ignoring statistical dependence and autocorrelation between their coefficients, which are prevalent when operating on time-series. To address this limitation, the Temporal Monge Alignment (TMA) [13, 14] was introduced as a pre-processing step to align temporal correlations by leveraging the Power Spectral density (PSD) of multivariate signals using Monge Optimal Transport mapping.
@@
Consider a multivariate signal $\mathbf{X}\triangleq[\mathbf{x}_{1},\dots,\mathbf{x}_{c}]^{\top}\in\mathbb{R}^{c\times\ell}$ of sufficient length. A standard assumption is that this signal follows a centered Gaussian distribution where sensors are uncorrelated and signals are periodic. This periodicity and uncorrelation structure implies that the signal’s covariance matrix is block diagonal, with each block having a circulant structure. A fundamental property of symmetric positive definite circulant matrices is their diagonalization [29] with real and positive eigenvalues in the Fourier basis $\mathbf{F}_{\ell}\in\mathbb{C}^{\ell\times\ell}$ of elements
@@
The Welch estimator [30] computes the PSD of a signal by averaging the squared Fourier transform of overlapping segments of the signal. Hence, the realization of the signal $\mathbf{X}\in\mathbb{R}^{c\times\ell}$ is decimated into overlapping segments $\{\mathbf{X}^{(1)},\dots,\mathbf{X}^{(L)}\}\subset\mathbb{R}^{c\times F}$ to estimate the PSD. The Welch estimator is defined as
@@
Let $\mathcal{N}(\mathbf{0},{\boldsymbol{\Sigma}}^{(s)})$ and $\mathcal{N}(\mathbf{0},{\boldsymbol{\Sigma}}^{(t)})$ be source and target centered Gaussian distributions respectively with covariance matrices following the structure (2) and PSDs denoted by $\mathbf{P}^{(s)}$ and $\mathbf{P}^{(t)}\in\mathbb{R}^{c\times F}$. Given a signal $\mathbf{X}\in\mathbb{R}^{c\times\ell}$ such that $\operatorname{vec}(\mathbf{X})\sim\mathcal{N}(\mathbf{0},{\boldsymbol{\Sigma}}^{(s)})$, the $F$-Monge mapping as defined by [13, 14] is
@@
For Gaussian distributions admitting the decomposition (2), the Wasserstein barycenter [31] admits an elegant closed-form solution. Consider $K$ centered Gaussian distributions admitting the decomposition (2) of PSDs $\mathbf{P}^{(1)},\dots,\mathbf{P}^{(K)}$. Their barycenter is also a centered Gaussian distribution $\mathcal{N}(\mathbf{0},\overline{{\boldsymbol{\Sigma}}})$ admitting the decomposition (2) with PSD
@@ 3 PSDNorm Layer
PSDNorm as a Drop-in Replacement
As discussed in Section 2, BatchNorm and InstanceNorm layers are commonly applied after each non-linearity in deep learning models, such as U-Time, to normalize feature maps. In contrast, PSDNorm offers a novel approach by aligning the PSD of feature maps to a barycenter PSD, providing a powerful alternative to classical normalization layers. Designed as a drop-in replacement, PSDNorm is optimized for modern hardware accelerators like GPUs, ensuring efficient execution. Once the deep-learning model is trained, PSDNorm operates as a test-time domain adaptation technique, allowing it to adapt to new data without additional training or access to training data. We define the normalized feature map as $\widetilde{\mathbf{G}}\triangleq\text{{PSDNorm}}(\mathbf{G})$. The following sections introduce the core components of PSDNorm and its implementation.
@@ 3.1 Core Components of the layer
In the following, we formally define PSDNorm and present each of its three main components: 1) PSD estimation, 2) running Riemannian barycenter estimation, and 3) $F$-Monge mapping computation. Given a batch =$\mathcal{B}=\{\mathbf{G}^{(1)},\dots,\mathbf{G}^{(N)}\}$ of $N$ pre-normalization feature maps, PSDNorm outputs a normalized batch ~=$\widetilde{\mathcal{B}}=\{\widetilde{\mathbf{G}}^{(1)},\dots,\widetilde{\mathbf{G}}^{(N)}\}$ with normalized PSD. Those three steps are detailed in the following and illustrated in the right part of Figure 1.
@@
First, the estimation of the PSD of each feature map is performed using the Welch method. The per-channel mean $\widehat{{\boldsymbol{\mu}}}^{(j)}$ is computed for each feature map $\mathbf{G}^{(j)}$ as $\widehat{{\boldsymbol{\mu}}}^{(j)}\triangleq\frac{1}{\ell}\sum_{l=1}^{\ell}\left[\mathbf{G}^{(j)}\right]_{:,l}\in\mathbb{R}^{c}\;.$ Then, the PSD of the centered feature map $\mathbf{G}^{(j)}-\widehat{{\boldsymbol{\mu}}}^{(j)}\mathbf{1}_{\ell}^{\top}$, denoted $\widehat{\mathbf{P}}^{(j)}$, is estimated as described in Equation (3). This centering step is required as feature maps are typically non-centered due to activation functions and convolution biases but they are assumed to have a stationary mean. The Welch estimation involves segmenting the centered feature map into overlapping windows, computing the Fourier transform of each window and then averaging them.
@@
Figure 2: Description of the running Riemanian barycenter. The barycenter of the batch $\widehat{\overline{\mathbf{P}}}_{\mathcal{B}}$ is estimated from the PSD of each batch sample. Then the running Riemanian barycenter is updated through an exponential average along the geodesic (- -), parameterized by $\alpha\in[0,1]$.
The PSDNorm aligns the PSD of each feature map to a barycenter PSD. This barycenter is computed during training by interpolating between the batch Wasserstein barycenter and the current running Riemanian barycenter using the geodesic associated with the Bures metric [32]. The batch barycenter is first computed from the current batch PSDs $\big{\{}\widehat{\mathbf{P}}^{(1)},\dots,\widehat{\mathbf{P}}^{(N)}\big{\}}$ using Equation 5. To ensure gradual adaptation, the running barycenter is updated via an exponential geodesic average with $\alpha\in[0,1]$:

### v3

# PSDNorm: Temporal Normalization for Deep Learning in Sleep Staging

@@ Abstract
Distribution shift poses a significant challenge in machine learning, particularly in biomedical applications using data collected across different subjects, institutions, and recording devices, such as sleep data. While existing normalization layers, BatchNorm, LayerNorm and InstanceNorm, help mitigate distribution shifts, when applied over the time dimension they ignore the dependencies and auto-correlation inherent to the vector coefficients they normalize. In this paper, we propose PSDNorm that leverages Monge mapping and temporal context to normalize feature maps in deep learning models for signals. Evaluations with architectures based on U-Net or transformer backbones trained on 10K subjects across 10 datasets, show that PSDNorm achieves state-of-the-art performance on unseen left-out datasets while being more robust to data scarcity.
@@ 1 Introduction
Machine learning techniques have achieved remarkable success in various domains, including computer vision, biology, audio processing, and language understanding. However, these methods face significant challenges when there are distribution shifts between training and evaluation datasets (Moreno-Torres et al., 2012). For example, in biological data, such as electroencephalography (EEG) signals, the distribution of the data can vary significantly. Indeed, data is collected from different subjects, electrode positions, and recording conditions. This paper focuses on sleep staging, a clinical task that consists in classifying periods of sleep in different stages based on EEG signals (Stevens and Clark, 2004). Depending on the dataset, the cohort can be composed of different age groups, sex repartition, health conditions, and recording conditions (O’Reilly et al., 2014; Quan et al., 1998; Marcus et al., 2013). Such variability brings shift in the distribution making it challenging for the model to generalize to unseen datasets.
@@
Normalization layers are widely used in deep learning to improve training stability and generalization. Common layers include BatchNorm (Ioffe and Szegedy, 2015), LayerNorm (Ba et al., 2016), and InstanceNorm (Ulyanov, 2016), which respectively compute statistics across the batch, normalize across all features within each sample, and normalize each channel independently within a sample. Some normalization methods target specific tasks, such as EEG covariance matrices (Kobler et al., 2022) or time-series forecasting (Kim et al., 2021), but they do not fully address spectral distribution shifts reflected in the temporal auto-correlations of signals. Other papers have proposed to adapt layer statistics to new domains (Li et al., 2016; Chang et al., 2019). In sleep staging, a simple normalization is often applied as preprocessing, e.g., standardizing signals over entire nights (Apicella et al., 2023) or short temporal windows (Chambon et al., 2018). Recent studies (Gnassounou et al., 2023; 2024) highlight the importance of considering temporal correlation and spectral content in normalization, proposing Temporal Monge Alignment (TMA), which aligns Power Spectral Density (PSD) to a common reference using Monge mapping, going beyond simple z-score normalization. However, these methods remain preprocessing steps that cannot be inserted as layers in the network architecture as it is done with BatchNorm, LayerNorm or InstanceNorm.
Deep Learning for Sleep Staging
@@
Sleep staging has been addressed by various neural network architectures, which process raw signals (Chambon et al., 2018; Perslev et al., 2021; Guillot and Thorey, 2021), spectrograms (Phan et al., 2023; 2019), or both (Phan et al., 2022a). More recent approaches involve transformer-based models that handle multimodal (Wang et al., 2024), spectrogram (Phan et al., 2022b), or heterogeneous inputs (Guo et al., 2024), offering improved modeling of temporal dependencies. However, most existing models are trained on relatively small cohorts, typically consisting of only a few hundred subjects, which limits their ability to generalize to diverse clinical settings. Notable exceptions include U-Sleep (Perslev et al., 2021), which was trained on a large-scale dataset and incorporates BatchNorm layers to mitigate data variability, and foundational models (Thapa et al., 2025; Fox et al., 2025; Deng et al., 2025) that achieve strong generalization from vast amount of data but require significant computational resources and are challenging to adapt without fine-tuning. Our focus is on developing smaller, efficient models that balance good generalization with ease of training and deployment in clinical practice.
@@
Figure 1: Description of normalization layers. The input shape is $(N,c,\ell)$ with batch size NN, channels $c$, and signal length $\ell$. BatchNorm estimates the mean $\widehat{\mu}$ and variance $\widehat{\sigma}^{2}$ over batch and time, and learns parameters $(\gamma,\beta)$ to normalize the input. PSDNorm estimates PSDs $\widehat{\mathbf{P}}$ over time and accounts for local temporal correlations. It computes the barycenter PSD $\widehat{\overline{\mathbf{P}}}$, updates it via a running Riemannian barycenter (6), and applies the filter $\widehat{\mathbf{H}}$ to normalize the input. The hyperparameter $f$ controls the extent of temporal correlation considered, thereby adjusting the strength of the normalization.
@@
In this work, we introduce the PSDNorm deep learning layer, a novel approach to address distribution shifts in machine learning for signals. PSDNorm leverages Monge Mapping to incorporate temporal context and normalize feature maps effectively. This layer enhances model robustness to new subjects at inference time. Unlike standard normalization layers such as LayerNorm or InstanceNorm, PSDNorm leverages the sequential nature of intermediate feature maps, as illustrated in Figure 1. We evaluate PSDNorm through extensive experiments on 10 sleep datasets. This evaluation covers 10M of samples across 10K subjects, using a leave-one-dataset-out (LODO) protocol with 3 different random seeds. To the best of our knowledge, such a large-scale and systematic evaluation has never been conducted before. PSDNorm achieves state-of-the-art performance and requires 4 times fewer labeled data to match the accuracy of the best baseline. Results highlight the potential of PSDNorm as a practical and efficient solution for tackling domain shifts in signals.
The paper is structured as follows: Section˜2 discusses existing normalization layers and pre-processing. Section˜3 introduces PSDNorm, followed by numerical results in Section˜4.
@@ 2 Related Works
In this section, we first review classical architectures for sleep staging and fundamental concepts of normalization layers. Then, we recall the Temporal Monge Alignment (TMA) method (Gnassounou et al., 2023) that aligns the PSD of signals using optimal transport.
Deep Learning for Sleep Staging
Numerous neural network architectures have been proposed for sleep staging, processing data in different formats as introduced in Section˜1. Different types of architectures have been explored, such as convolutional neural networks (CNNs) (Chambon et al., 2018), recurrent neural networks (RNNs) (Supratak et al., 2017; Phan et al., 2019), and more recently transformers (Phan et al., 2022b; Wang et al., 2024; Guo et al., 2024), which have shown promise in modeling temporal dependencies in sleep data. While many models are typically evaluated on a limited number of datasets, the work by (Perslev et al., 2021) introduced U-Sleep, a model trained on a large-scale dataset of sleep recordings. Their architecture, based on U-Time (Perslev et al., 2019), incorporates BatchNorm layers to mitigate data variability, and they employ a domain generalization approach: training a single model on a sufficiently diverse set of domains to ensure it generalizes to unseen datasets without additional adaptation. This architecture is composed of encoder-decoder blocks with skip connections, allowing the model to capture both local and global features of the sleep signals effectively. Each encoder and decoder block consists of convolutional layers followed by BatchNorm and non-linear activation functions, enabling the model to learn robust representations of the input data. A more detailed description of U-Time is provided in Section˜A.3.
@@
Normalization layers enhance training and robustness in deep neural networks. The most common are BatchNorm (Ioffe and Szegedy, 2015), InstanceNorm (Ulyanov, 2016), and LayerNorm (Ba et al., 2016). BatchNorm normalizes feature maps using batch and time statistics, ensuring zero mean and unit variance. The output is adjusted with learnable parameters. InstanceNorm normalizes each channel per sample using its own statistics, independent of the batch (see Fig. 1). Popular in time-series forecasting, it is used in RevIN (Kim et al., 2021), which reverses normalization after decoding. LayerNorm normalizes across all channels and time steps within each sample, with learnable scaling and shifting. While these normalization layers are widely employed, they operate on vectors ignoring statistical dependence and autocorrelation between their coefficients, which are prevalent when operating on time-series. To address this limitation, the Temporal Monge Alignment (TMA) (Gnassounou et al., 2023; 2024) was introduced as a pre-processing step to align temporal correlations by leveraging the Power Spectral density (PSD) of multivariate signals using Monge Optimal Transport mapping.
@@
Consider a multivariate signal $\mathbf{X}\triangleq[\mathbf{x}_{1},\dots,\mathbf{x}_{c}]^{\top}\in\mathbb{R}^{c\times\ell}$ of sufficient length. A standard assumption is that this signal follows a centered Gaussian distribution where sensors are uncorrelated and signals are periodic. This periodicity and uncorrelation structure implies that the signal’s covariance matrix is block diagonal, with each block having a circulant structure. A fundamental property of symmetric positive definite circulant matrices is their diagonalization (Gray, 2006) with real and positive eigenvalues in the Fourier basis $\mathbf{F}_{\ell}\in\mathbb{C}^{\ell\times\ell}$ of elements
@@
The Welch estimator (Welch, 1967) computes the PSD of a signal by averaging the squared Fourier transform of overlapping segments of the signal. Hence, the realization of the signal $\mathbf{X}\in\mathbb{R}^{c\times\ell}$ is decimated into overlapping segments $\{\mathbf{X}^{(1)},\dots,\mathbf{X}^{(L)}\}\subset\mathbb{R}^{c\times f}$ to estimate the PSD. The Welch estimator is defined as
@@
Let $\mathcal{N}(\mathbf{0},{\boldsymbol{\Sigma}}^{(s)})$ and $\mathcal{N}(\mathbf{0},{\boldsymbol{\Sigma}}^{(t)})$ be source and target centered Gaussian distributions respectively with covariance matrices following the structure (2) and PSDs denoted by $\mathbf{P}^{(s)}$ and $\mathbf{P}^{(t)}\in\mathbb{R}^{c\times f}$. Given a signal $\mathbf{X}\in\mathbb{R}^{c\times\ell}$ such that $\operatorname{vec}(\mathbf{X})\sim\mathcal{N}(\mathbf{0},{\boldsymbol{\Sigma}}^{(s)})$, the $f$-Monge mapping as defined by (Gnassounou et al., 2023; 2024) is
@@
For Gaussian distributions admitting the decomposition (2), the Wasserstein barycenter (Agueh and Carlier, 2011) admits an elegant closed-form solution. Consider $K$ centered Gaussian distributions admitting the decomposition (2) of PSDs $\mathbf{P}^{(1)},\dots,\mathbf{P}^{(K)}$. Their barycenter is also a centered Gaussian distribution $\mathcal{N}(\mathbf{0},\overline{{\boldsymbol{\Sigma}}})$ admitting the decomposition (2) with PSD
@@ 3 PSDNorm Layer
PSDNorm is a novel normalization layer that can be used as a drop-in replacement for layers like BatchNorm or InstanceNorm. Instead of simple standardization, it aligns the Power Spectral Density (PSD) of feature maps to a running barycenter PSD. This approach, optimized for modern hardware, enhances model robustness to new subjects at inference time without retraining. We define the normalized feature map as $\widetilde{\mathbf{G}}\triangleq\text{{PSDNorm}}(\mathbf{G})$. The following sections introduce the core components of PSDNorm and its implementation.
@@ 3.1 Core Components of the layer
In the following, we formally define PSDNorm and present each of its three main components: 1) PSD estimation, 2) running Riemannian barycenter estimation, and 3) $f$-Monge mapping computation. Given a batch $\mathcal{B}=\{\mathbf{G}^{(1)},\dots,\mathbf{G}^{(N)}\}$ of $N$ pre-normalization feature maps, PSDNorm outputs a normalized batch $\widetilde{\mathcal{B}}=\{\widetilde{\mathbf{G}}^{(1)},\dots,\widetilde{\mathbf{G}}^{(N)}\}$ with normalized PSD. Those three steps are detailed in the following and illustrated in the right part of Figure 1.
@@
Figure 2: Description of the running Riemanian barycenter. The barycenter of the batch $\widehat{\overline{\mathbf{P}}}_{\mathcal{B}}$ is estimated from the PSD of each batch sample.
First, the estimation of the PSD of each feature map is performed using the Welch method. The per-channel mean $\widehat{{\boldsymbol{\mu}}}^{(j)}$ is computed for each feature map $\mathbf{G}^{(j)}$ as $\widehat{{\boldsymbol{\mu}}}^{(j)}\triangleq\frac{1}{\ell}\sum_{l=1}^{\ell}\left[\mathbf{G}^{(j)}\right]_{:,l}\in\mathbb{R}^{c}\;.$
Then, the PSD of the centered feature map $\mathbf{G}^{(j)}-\widehat{{\boldsymbol{\mu}}}^{(j)}\mathbf{1}_{\ell}^{\top}$, denoted $\widehat{\mathbf{P}}^{(j)}$, is estimated as described in Equation (3). This centering step is required as feature maps are typically non-centered due to activation functions and convolution biases but they are assumed to have a stationary mean. The Welch estimation involves segmenting the centered feature map into overlapping windows, computing the Fourier transform of each window and then averaging them.
@@
The PSDNorm aligns the PSD of each feature map to a barycenter PSD. This barycenter is computed during training by interpolating between the batch Wasserstein barycenter and the current running Riemanian barycenter using the geodesic associated with the Bures metric (Bhatia et al., 2019). The batch barycenter is first computed from the current batch PSDs $\big\{\widehat{\mathbf{P}}^{(1)},\dots,\widehat{\mathbf{P}}^{(N)}\big\}$ using Equation˜5. To ensure gradual adaptation, the running barycenter is updated via an exponential geodesic average with $\alpha\in[0,1]$: